Circular No. 6901

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bw IN ESO 184-G82 E. M. Sadler, University of Sydney; R. A. Stathakis and B. J. Boyle,

Anglo-Australian Observatory; and R. D. Ekers, Australia Telescope National Facility, report: “The spectrum of the variable object in ESO 184-G82 (IAUC 6895, 6896, 6899) closely resembles the pre-maximum spectrum of the type-Ib supernova 1983N (Richtler and Sadler 1983, A.Ap. 128, L3). The detection of prompt radio emission (IAUC 6896; see http://www.

narrabri.atnf.csiro.au/public/grb for more recent data) is also typi-

cal of type-Ib supernovae (though this object is ∼ 30 times more luminous at radio wavelengths than was SN 1983N). It therefore seems certain that the object in ESO 184-G82 is a supernova discovered before maximum light. The optical lightcurve (IAUC 6895, 6899) implies that the supernova explosion probably occurred sometime during Apr. 21–27 — i.e., consistent with the supernova and the γ-ray burst GRB 980425 occurring simultaneously. We suggest that the two events are associated, and that we may be seeing a supernova in which the core has collapsed to a black hole (plus accretion disk) rather than a neutron star, after models proposed by Woosley (1993, Ap.J. 405, 273) and Paczynski (1998, Ap.J. 484, L45).”

OGLE-2 SYSTEM FOR MICROLENSING EVENTS A. Udalski and M. Szymanski, on behalf of the OGLE collaboration,

announce: “The Early Warning System (EWS) designed for the detection of microlensing events in progress and successfully implemented during the first phase of the OGLE experiment (cf. IAUC 5997, 6164, 6177) has been reinstalled for the second phase of the experiment. About 25 million stars are currently monitored in the Galactic bulge each night with the 1.3-m Warsaw telescope at Las Campanas Observatory. The expected rate of microlensing events is ∼ 50 per observing season (based on preliminary results from the 1997 season). So far in the 1998 season, 16 microlensing event candidates have been detected, five of them still in their early phases before or shortly after maximum light. We direct attention to the event OGLE-1998-BUL-14, which is expected to reach high amplification with maximum around May 20 and be brighter than I = 15. Finding charts, photometry, and other information about

OGLE events can be found at http://www.astrouw.edu.pl/~ftp/ogle
or ftp://sirius.astrouw.edu.pl/ogle/ogle2/ews. Astronomers inter-

ested in follow-up observations and who wish to receive e-mail notification

should send a request to ogle-ews@sirius.astrouw.edu.pl.”

1998 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green